# STAR: A Simple Training-free Approach for Recommendations using Large Language Models

### v1

@@ Abstract.
Recent progress in large language models (LLMs) offers promising new approaches for recommendation system (RecSys) tasks. While the current state-of-the-art methods rely on fine-tuning LLMs to achieve optimal results, this process is costly and introduces significant engineering complexities. Conversely, methods that bypass fine-tuning and use LLMs directly are less resource-intensive but often fail to fully capture both semantic and collaborative information, resulting in sub-optimal performance compared to their fine-tuned counterparts. In this paper, we propose a Simple Training-free Approach for Recommendation (STAR), a framework that utilizes LLMs and can be applied to various recommendation tasks without the need for fine-tuning. Our approach involves a retrieval stage that uses semantic embeddings from LLMs combined with collaborative user information to retrieve candidate items. We then apply an LLM for pairwise ranking to enhance next-item prediction. Experimental results on the Amazon Review dataset show competitive performance for next item prediction, even with our retrieval stage alone. Our full method achieves Hits@10 performance of +23.8% on Beauty, +37.5% on Toys and Games, and -1.8% on Sports and Outdoors relative to the best supervised models. This framework offers an effective alternative to traditional supervised models, highlighting the potential of LLMs in recommendation systems without extensive training or custom architectures.
Recommender System, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Information systems Recommender systems
1. Introduction
Personalized recommendation systems have become indispensable tools for enhancing user experiences and driving engagement across a wide range of online platforms. Recent advances in large language models (LLMs) present new opportunities for addressing recommendation tasks (brown2020language, ; team2023gemini, ; lin2023can, ; zhao2023recommender, ; li2023large, ; chen2024large, ; tsai-etal-2024-leveraging, ; wu2024survey, ). Current strategies primarily involve utilizing LLMs as either feature encoders (sun2019bert4rec, ; Ding2022, ; hou2022towards, ; hou2023learning, ; singh2023better, ; yuan2023go, ; harte2023leveraging, ; gong2023unified, ; li2023text, ; liu2024once, ; li2024enhancing, ; ren2024representation, ; rajput2024recommender, ; zheng2024adapting, ; sheng2024language, ; hu2024enhancing, ) or as scoring and ranking functions (wang2023zero, ; wang2023drdt, ; hou2024large, ; wang-etal-2024-recmind, ; xu2024prompting, ; zhao2024let, ; liang2024taxonomy, ). When LLMs are employed as feature encoders, there is potential for transfer learning and cross-domain generalization by initializing embedding layers with LLM embeddings, although this approach requires extensive training. On the other hand, using LLMs for scoring and ranking demonstrates the ability to leverage their reasoning capabilities to address recommendation tasks. However, these models still lag behind the performance of fine-tuned models due to a lack of collaborative knowledge.
Figure 1. Retrieval pipeline uses scoring rules that combine semantic and collaborative information with temporal, weight, and rating factors to score unseen items without requiring any fine-tuning.
@@ 1. Introduction
The primary motivation of this work is to develop a general framework that serves as a generalist across multiple recommendation domains. We demonstrate that recent advancements in LLMs align with this vision, effectively functioning as generalists without requiring any domain-specific fine-tuning. Based on our findings, we present a Simple Training-free Approach for Recommendation (STAR) framework using LLMs. The STAR framework involves two stages: Retrieval and Ranking. The Retrieval stage scores new items using a combination of semantic similarity and collaborative commonality to the items in a user’s history. Here, we utilize LLM-based embeddings to determine semantic similarity. Additionally, a temporal factor gives priority to user’s recent interactions, and a rating factor aligns with user preferences to rank items within a specific set (See Figure 1 and Section 3.2). The Ranking stage leverages the reasoning capabilities of LLMs to adjust the rankings of the initially retrieved candidates. Specifically, we assess various LLM-based ranking approaches, including point-wise, pair-wise, and list-wise methods, while also determining the key information needed for the LLM to better understand user preferences and make accurate predictions (Section 3.3). Our experimental evaluation shows competitive performance across a diverse range of recommendation datasets, all without the need for supervised training or the development of custom-designed architectures.
We present extensive experimental results on the Amazon Review dataset (mcauley2015image, ; he2016ups, ). Our findings are as follow: (1) Our retrieval pipeline, comprised of both semantic relationship and collaborative information, demonstrates competitive results compared to a wide range of fine-tuned methods. LLM embeddings allow for an effective method to calculate semantic similarity; (2) We show that pair-wise ranking further improves upon our retrieval performance, while point-wise and list-wise methods struggle to achieve similar improvements; and (3) We illustrate that collaborative information is a critical component that adds additional benefits to the semantic information throughout our system, in both the retrieval and ranking stages.
2. Related Works
LLM as a Feature Encoder for RecSys.
@@ 2. Related Works
Recommendation systems typically leverage feature encoders to transform item and user profiles into suitable representations for model training. Traditionally, ID-based systems relied on one-hot encoding for structured features (zhou2018deep, ; wang2021dcn, ). However, recent advancements in LLMs have enabled the utilization of text encoders to capture rich semantic information from item metadata and user profiles (reimers2019sentence, ; cer-etal-2018-universal, ; ni2021sentence, ; lee2024gecko, ) To further optimize these representations for specific applications, researchers have explored several approaches: (1) mapping continuous LLM embeddings into discrete tokens using vector quantization and training a subsequent generative model (hou2023learning, ; singh2023better, ; rajput2024recommender, ; zheng2024adapting, ); (2) training sequential models by initializing the embedding layer with LLM embeddings (sun2019bert4rec, ; yuan2023go, ; hu2024enhancing, ); and (3) training models to directly compute the relevance between item and user embeddings (i.e., embeddings of user selected items) (Ding2022, ; hou2022towards, ; gong2023unified, ; li2023text, ; liu2024once, ; li2024enhancing, ; ren2024representation, ; sheng2024language, ). While optimizing representations can improve recommendation performance, this often comes at the cost of increased training expenses and reduced generalizability. In this work, we demonstrate that LLM embeddings can be directly used as effective item representations, yielding strong results in sequential recommendation tasks without requiring extensive optimization. This finding aligns with those of (harte2023leveraging, ), but differs by usage of novel scoring rules that incorporates collaborative and temporal information.
LLM as a Scoring and Ranking function for RecSys.
Recent studies show that LLMs can recommend items by understanding user preferences or past interactions in natural language. This is achieved through generative selection prompting, where the model ranks and selects top recommended items from a set of candidates (wang2023zero, ; wang2023drdt, ; hou2024large, ; wang-etal-2024-recmind, ; xu2024prompting, ; zhao2024let, ; liang2024taxonomy, ). However, these studies show that LLMs alone are less effective than models fine-tuned on user-item interaction data, which leverage collaborative knowledge. To bridge the gap between collaborative knowledge and the semantic understanding of LLMs, recent efforts have focused on fine-tuning the models with interaction data, though this approach is also costly (geng2022recommendation, ; zhang2023recommendation, ; bao2023tallrec, ; xu2024openp5, ; tan2024idgenrec, ; kim2024large, ).
LLM as a Ranker for Information Retrieval.
Ranking using LLMs has been widely adopted in document retrieval (zhu2023large, ; wang2024large, ). Recent studies indicate that LLMs surpass traditional supervised cross-encoders in zero-shot passage ranking. Three prompting approaches have emerged: (1) point-wise: LLMs directly evaluate relevance using numerical scores or binary judgments (liang2022holistic, ; zhuang2023beyond, ), but this method struggles with capturing the relative importance of passages; (2) pair-wise: LLMs express preferences between item pairs, which is effective but inefficient due to the high number of calls required (qin-etal-2024-large, ); (3) list-wise: LLMs compare multiple passages simultaneously (sun2023chatgpt, ), but performance heavily relies on the model’s semantic prior and reasoning capabilities (qin-etal-2024-large, ). In this paper, we investigate the potential of LLM ranking for recommendation tasks, which are subjective in nature, unlike the deterministic nature of document retrieval.
3. STAR: Simple Training-Free Approach
@@
Figure 2. STAR Framework overview. We use the semantic relationship scores in $R_{\text{S}}$ and the collaborative relationship scores in $R_{\text{C}}$ to score the items in the user history compared to new items to recommend. The final score for one new item is a weighted average from the semantic relationship and collaborative relationship scores, with additional weights from the user’s ratings $r$ and a temporal decay $\lambda<1$ which prioritize recent interactions. The top scoring retrieved items are sent to the LLM Ranking, where we can use point-wise, pair-wise, or list-wise ranking approaches to further improve upon the scoring of recommended items.
@@
This section initially outlines the problem formulation (Section 3.1). Subsequently, we detail the proposed retrieval (Section 3.2) and ranking pipelines (Section 3.3).
3.1. Sequential Recommendation
@@
The sequential recommendation task aims to predict the next item a user will interact with based on their interaction history. For a user $u\in U$, where $U$ is the set of all users, the interaction history is represented as a sequence of items $S_{u}=\{s_{1},s_{2},\dots,s_{n}\}$, with each $s_{i}\in I$ belonging to the set of all items $I$. Each user history item $s_{i}$ is associated with a rating $r_{i}\in\{1,2,3,4,5\}$ given by the user $u$. The goal is to predict the next item $s_{n+1}\in I$ that the user is most likely to interact with.
3.2. Retrieval Pipeline
@@
Understanding how similar a candidate item is to the items in a user’s interaction history $s_{i}\in S_{u}$ is key to accurately gauging how well candidate items align with user preferences. Here we leverage LLM embedding models, where we pass in custom text prompts representing items and collect embedding vectors of dimension $d_{e}$. We construct a prompt based on the item information and metadata, including the title, description, category, brand, sales ranking, and price. We omit metadata fields like Item ID and URL, as those fields contain strings that can contain spurious lexical similarity (e.g., IDs: “000012”, “000013’ or URLs: “https://abc.com/uxrl”, “https://abc.com/uxrb”) and can reduce the uniformity of the embedding space and make it difficult to distinguish between semantically different items (See Appendix A.1 for the full prompt). We collect embeddings for each item $i\in I$, resulting in $E\in\mathbb{R}^{n\times d_{e}}$, where $n$ is number of items in $I$.
The semantic relationship between two items ($i_{a},i_{b}$) is then calculated using the cosine similarity between their embeddings $E_{i_{a}},E_{i_{b}}\in E$. This measure provides a numerical representation of how closely related the items are in semantic space. For our experiments, we precompute the entire semantic relationship matrix $R_{\text{S}}\in\mathbb{R}^{n\times n}$. For many domains, this is a practical solution. However, if $|I|$ is very large, Approximate Nearest Neighbor methods (guo2020accelerating, ; sun2024soar, ) are efficient approaches to maintain quality and reduce computation.
@@
Semantic similarity between a candidate item and items in a user’s interaction history is a helpful cue for assessing the similarity of items based on the item information. However, this alone does not fully capture the engagement interactions of items by multiple users. To better understand the collaborative relationship, we consider how frequently different combinations of items are interacted with by users. These shared interaction patterns can provide strong indicators of how likely the candidate item is to resonate with a broader audience with similar preferences. For each item $i\in I$, we derive an interaction array that represents user interactions, forming a set of sparse user-item interaction arrays $C\in\mathbb{R}^{n\times m}$, where $m$ is number of users in $U$. The collaborative relationship between two items ($i_{a},i_{b}$) is then computed by using the cosine similarity between their sparse arrays $C_{i_{a}},C_{i_{b}}\in C$, capturing the normalized co-occurrence of the items. To streamline the process, we pre-compute and store these values in a collaborative relationship matrix $R_{\text{C}}\in\mathbb{R}^{n\times n}$, which is typically very sparse.
@@
(1) $\texttt{score}(x)=\frac{1}{n}\sum_{j=1}^{n}r_{j}\cdot\lambda^{t_{j}}\cdot\left[a\cdot R_{\text{S}}^{xj}+(1-a)\cdot R_{\text{C}}^{xj}\right]$
@@
3.3. Ranking Pipeline
Figure 3. Prompt overview for the ranking pipeline. The prompt includes history items, candidate items, and instructions for the ranking strategy. Each item is represented by metadata, along with additional details such as popularity and co-occurrence, formatted in JSON. Full prompt is available in Appendix A.2.
@@ 3.3. Ranking Pipeline
After retrieving the top $k$ items, denoted as $I_{k}$, from the initial retrieval process, a LLM is employed to further rank these items to enhance the overall ranking quality. The items in $I_{k}$ are already ordered based on scores from the retrieval framework, which reflect semantic, collaborative, and temporal information. We intentionally incorporate this initial order into the ranking process to enhance both efficiency and effectiveness. This framework then leverages the capabilities of the LLM to better capture user preference, complex relationships and contextual relevance among the items.
3.3.1. Rank schema

### v2

@@ Abstract
Recent progress in large language models (LLMs) offers promising new approaches for recommendation system tasks. While the current state-of-the-art methods rely on fine-tuning LLMs to achieve optimal results, this process is costly and introduces significant engineering complexities. Conversely, methods that directly use LLMs without additional fine-tuning result in a large drop in recommendation quality, often due to the inability to capture collaborative information. In this paper, we propose a Simple Training-free Approach for Recommendation (STAR), a framework that utilizes LLMs and can be applied to various recommendation tasks without the need for fine-tuning, while maintaining high quality recommendation performance. Our approach involves a retrieval stage that uses semantic embeddings from LLMs combined with collaborative user information to retrieve candidate items. We then apply an LLM for pairwise ranking to enhance next-item prediction. Experimental results on the Amazon Review dataset show competitive performance for next item prediction, even with our retrieval stage alone. Our full method achieves Hits@10 performance of +23.8% on Beauty, +37.5% on Toys & Games, and -1.8% on Sports & Outdoors relative to the best supervised models. This framework offers an effective alternative to traditional supervised models, highlighting the potential of LLMs in recommendation systems without extensive training or custom architectures.
STAR: A Simple Training-free Approach for
Recommendations using Large Language Models
Dong-Ho Lee1thanks:   Authors contributed equally.thanks:   Work done during internship at Google DeepMind.,  Adam Kraft2,  Long Jin3,  Nikhil Mehta2, Taibai Xu3 , Lichan Hong2Ed H. Chi2Xinyang Yi2, 1Information Sciences Institute, University of Southern California 2Google DeepMind  3Google {dongho.lee}@usc.edu, {adamkraft,longjin,nikhilmehta,taibaixu,lichan,edchi,xinyang}@google.com
1 Introduction
Figure 1: A Motivating Example. LLMs can be utilized in RecSys through (a) prompting, (b) fine-tuning on user-item interactions, and (c) using LLMs as feature encoders for training subsequent models. However, (a) cannot leverage collaborative knowledge, while (b) and (c) require extensive training and large-scale interaction data. Our framework STAR integrates collaborative knowledge into LLMs without additional training.
Recent advances in large language models (LLMs) present new opportunities for addressing recommendation tasks (Brown et al., 2020; Team et al., 2023; Lin et al., 2023; Zhao et al., 2023; Li et al., 2023b; Chen et al., 2024; Tsai et al., 2024; Wu et al., 2024). Current approaches primarily leverage LLMs in three ways: (1) advanced prompting Wang and Lim (2023); Wang et al. (2023); Hou et al. (2024); Wang et al. (2024b); Xu et al. (2024a); Zhao et al. (2024); Liang et al. (2024); (2) as base models for direct fine-tuning Geng et al. (2022); Zhang et al. (2023); Bao et al. (2023); Xu et al. (2024b); Tan et al. (2024); Kim et al. (2024); and (3) as feature encoders for fine-tuning subsequent generative models (Hou et al., 2023; Singh et al., 2023; Rajput et al., 2024; Zheng et al., 2024) or sequential models Sun et al. (2019); Yuan et al. (2023); Hu et al. (2024). While prompting LLMs for scoring and ranking utilizes their reasoning abilities, these models often largely underperform compared to fine-tuned approaches due to the absence of collaborative knowledge derived from user-item interactions. Conversely, using LLMs for fine-tuning, whether as base models or feature encoders, enhances model performance by leveraging their strong semantic understanding. However, this requires extensive training and large-scale interaction data.
@@ 1 Introduction
The main objective of this work is to integrate the semantic capabilities of LLMs with collaborative knowledge from user-item interaction data, all without requiring additional training. To achieve this, we present a Simple Training-free Approach for Recommendation (STAR) framework using LLMs. The STAR framework involves two stages: Retrieval and Ranking. The Retrieval stage scores new items using a combination of semantic similarity and collaborative commonality to the items in a user’s history. Here, we utilize LLM-based embeddings to determine semantic similarity. Additionally, a temporal factor gives priority to user’s recent interactions, and a rating factor aligns with user preferences to rank items within a specific set (§3.2). The Ranking stage leverages the reasoning capabilities of LLMs to adjust the rankings of the initially retrieved candidates. Specifically, we assess various LLM-based ranking approaches, including point-wise, pair-wise, and list-wise methods, while also determining the key information needed for the LLM to better understand user preferences and make accurate predictions (§3.3). Our experimental evaluation shows competitive performance across a diverse range of recommendation datasets, all without the need for supervised training or the development of custom-designed architectures.
We present extensive experimental results on the Amazon Review dataset (McAuley et al., 2015; He and McAuley, 2016). Our findings are as follow:
  1. 1.
    Our retrieval pipeline itself, comprised of both semantic relationship and collaborative information, achieves Hits@10 performance of +17.3% on Beauty, +26.2% on Toys & Games, and -5.5% on Sports & Outdoors relative to the best supervised models.
  2. 2.
    We show that pair-wise ranking further improves upon our retrieval performance, while point-wise and list-wise methods struggle to achieve similar improvements.
  3. 3.
    We illustrate that collaborative information is a critical component that adds additional benefits to the semantic information throughout our system, in both the retrieval and ranking stages.
These findings show that it is possible to build a recommendation system utilizing LLMs without additional fine-tuning that can significantly close the quality gap of fully fine-tuned systems, and in many cases even achieve higher quality.
2 Related Works
Recent studies have explored the role of LLMs in recommendation systems through three primary approaches: (1) using prompting for scoring and ranking, (2) fine-tuning as a base model, and (3) serving as a feature encoder.
LLM prompting for scoring and ranking.
LLMs can generate recommendations by understanding user preferences or past interactions expressed in natural language.
This is typically achieved through generative selection prompting, where the model ranks and selects top items from a given candidate set (Wang and Lim, 2023; Wang et al., 2023; Hou et al., 2024; Wang et al., 2024b; Xu et al., 2024a; Zhao et al., 2024; Liang et al., 2024). Another line of work applies ranking prompting (Dai et al., 2023), inspired by LLM-based ranking in information retrieval (Zhu et al., 2023; Wang et al., 2024a), using point-wise (Liang et al., 2022; Zhuang et al., 2023), pair-wise (Qin et al., 2024), or list-wise (Sun et al., 2023; Qin et al., 2024) approaches. However, LLM prompting alone, without a retrieval stage to pre-select candidate items for scoring and ranking, is less effective than fine-tuned models due to lack of collaborative knowledge derived from user-item interaction data. As a result, many approaches use a fine-tuned model for candidate retrieval and an LLM for ranking. However, in this setup, overall performance is primarily determined by the retrieval quality rather than the LLM itself.
LLM as base model for fine-tuning.
To integrate collaborative knowledge with the semantic understanding of LLMs, recent studies have explored fine-tuning using user-item interaction data. While this improves recommendation performance, it requires extensive training and large-scale interaction datasets (Geng et al., 2022; Zhang et al., 2023; Bao et al., 2023; Xu et al., 2024b; Tan et al., 2024; Kim et al., 2024).
LLM as feature encoder.
@@ 2 Related Works
LLMs can also be used as text encoders to capture richer semantic information from item metadata and user profiles Reimers and Gurevych (2019); Cer et al. (2018); Ni et al. (2021); Lee et al. (2024). To further optimize these representations, researchers have explored several approaches: (1) mapping continuous LLM embeddings into discrete tokens using vector quantization and training a subsequent generative model (Hou et al., 2023; Singh et al., 2023; Rajput et al., 2024; Zheng et al., 2024); (2) training sequential models by initializing the embedding layer with LLM embeddings (Sun et al., 2019; Yuan et al., 2023; Hu et al., 2024); and (3) training models to directly compute the relevance between item and user embeddings (i.e., embeddings of user selected items) (Ding et al., 2022; Hou et al., 2022; Gong et al., 2023; Li et al., 2023a; Liu et al., 2024; Li et al., 2024; Ren et al., 2024; Sheng et al., 2024).
In this work, we utilize LLMs as both feature encoders and ranking functions by integrating collaborative knowledge from user-item interaction data. Our findings show that LLM embeddings can serve as effective item representations, achieving strong results in sequential recommendation tasks without extensive optimization. This aligns with (Harte et al., 2023) but differs in our use of novel scoring rules that incorporate both collaborative and temporal information.
3 STAR: Simple Training-Free RecSys
@@
Figure 2: STAR Framework overview. We use the semantic relationship scores in $R_{\text{S}}$ and the collaborative relationship scores in $R_{\text{C}}$ to score the items in the user history compared to new items to recommend. The final score for one new item is a weighted average from the semantic relationship and collaborative relationship scores, with additional weights from the user’s ratings $r$ and a temporal decay $\lambda<1$ which prioritize recent interactions. The top scoring retrieved items are sent to the LLM Ranking, where we can use point-wise, pair-wise, or list-wise ranking approaches to further improve upon the scoring of recommended items.
@@
This section initially outlines the problem formulation (§3.1). Subsequently, we detail the proposed retrieval (§3.2) and ranking pipelines (§3.3).
3.1 Sequential Recommendation
@@
The sequential recommendation task predicts the next item a user will interact with based on their interaction history. For a user $u\in U$, where $U$ is the set of all users, the interaction history is represented as a sequence of items $S_{u}=\{s_{1},s_{2},\dots,s_{n}\}$, with each item $s_{i}\in I$ belonging to the set of all items $I$. Each user history item $s_{i}$ is associated with a rating $r_{i}\in\{1,2,3,4,5\}$ given by the user $u$. The goal is to predict the next item $s_{n+1}\in I$ that the user is most likely to interact with.
3.2 Retrieval Pipeline
@@
Understanding how similar a candidate item is to the items in a user’s interaction history $s_{i}\in S_{u}$ is key to accurately gauging how well candidate items align with user preferences. Here we leverage LLM embedding models, where we pass in text prompts representing items and collect embedding vectors of dimension $d_{e}$. We construct a prompt based on the item information and metadata, which can include fields like title, description, category, brand, sales ranking, price, etc. (See Appendix A.3 for the full prompt details). We collect embeddings for each item $i\in I$, resulting in $E\in\mathbb{R}^{n\times d_{e}}$, where $n$ is number of total items in $I$.
The semantic relationship between two items ($i_{a},i_{b}$) is then calculated using the cosine similarity between their embeddings $E_{i_{a}},E_{i_{b}}\in E$. This measure provides a numerical representation of how closely related the items are in semantic space. For our experiments, we precompute the entire semantic relationship matrix $R_{\text{S}}\in\mathbb{R}^{n\times n}$. For many domains, this is a practical solution. However, if $|I|$ is very large, Approximate Nearest Neighbor methods (Guo et al., 2020; Sun et al., 2024) are efficient approaches to maintain quality and reduce computation.
@@
Semantic similarity between a candidate item and items in a user’s interaction history is a helpful cue for assessing the similarity of items based on the item information. However, this alone does not fully capture the engagement interactions of items by multiple users. To better understand the collaborative relationship, we consider how frequently different combinations of items are interacted with by users. These shared interaction patterns can provide strong indicators of how likely the candidate item is to resonate with a broader audience with similar preferences. For each item $i\in I$, we derive an interaction array that represents user interactions, forming a set of sparse user-item interaction arrays $C\in\mathbb{R}^{n\times m}$, where $m$ is number of users in $U$. The collaborative relationship between two items ($i_{a},i_{b}$) is then computed by using the cosine similarity between their sparse arrays $C_{i_{a}},C_{i_{b}}\in C$, capturing the normalized co-occurrence of the items. To streamline the process, we pre-compute and store these values in a collaborative relationship matrix $R_{\text{C}}=\frac{C\cdot C^{\top}}{\|C\|\|C^{\top}\|}\in\mathbb{R}^{n\times n}$, which is typically very sparse.
@@
$\texttt{score}(x)=\frac{1}{n}\sum_{j=1}^{n}r_{j}\lambda^{t_{j}}\left[aR_{\text{S}}^{xj}+(1-a)R_{\text{C}}^{xj}\right]$ (1)
@@
3.3 Ranking Pipeline
@@
After retrieving the top $k$ items, denoted as $I_{k}$, from the initial retrieval process, a LLM is employed to further rank these items to enhance the overall next-item recommendation quality. The items in $I_{k}$ are already ordered based on scores from the retrieval framework, which reflect semantic, collaborative, and temporal information. We intentionally incorporate this initial order into the ranking process to enhance both efficiency and effectiveness. This framework then leverages the capabilities of the LLM to better capture user preference, complex relationships and contextual relevance among the items.
3.3.1 Rank schema